# REST API Testing in DevOps: A Study on an Evolving Healthcare IoT Application

### v1

@@ Abstract
Healthcare Internet of Things (IoT) applications often integrate various third-party healthcare applications and medical devices through REST APIs, resulting in complex and interdependent networks of REST APIs. Oslo City’s healthcare department collaborates with various industry partners to develop such healthcare IoT applications enriched with a diverse set of REST APIs. Following the DevOps process, these REST APIs continuously evolve to accommodate evolving needs such as new features, services, and devices. Oslo City’s primary goal is to utilize automated solutions for continuous testing of these REST APIs at each evolution stage, thereby ensuring their dependability. Although the literature offers various automated REST API testing tools, their effectiveness in regression testing of the evolving REST APIs of healthcare IoT applications within a DevOps context remains undetermined. This paper evaluates state-of-the-art and well-established REST API testing tools—specifically, RESTest, EvoMaster, Schemathesis, RESTler, and RestTestGen—for the regression testing of a real-world healthcare IoT application, considering failures, faults, coverage, regressions, and cost. We conducted experiments using all accessible REST APIs (17 APIs with 120 endpoints), and 14 releases evolved during DevOps. Overall, all tools generated tests leading to several failures, 18 potential faults, up to 84% coverage, 23 regressions, and over 80% cost overhead.
Keywords Black-box Testing, Fuzzing, REST API, Web Services, DevOps, CI/CD, Healthcare IoT
@@ 1 Introduction
Healthcare Internet of Things (IoT) applications commonly use a cloud-based architecture as a hub for stakeholders such as healthcare professionals, caretakers, and patients [1]. This involves integrating various healthcare applications serving different purposes (such as electronic health records (EHR) for managing patient data) and medical devices (like medicine dispensers) assigned to patients. These integrations into the IoT cloud are facilitated through Application Programming Interfaces (APIs) following the Representational State Transfer (REST) architecture—concisely REST APIs. A malfunction in REST APIs can significantly dismantle the healthcare IoT cloud, potentially making crucial healthcare services unavailable to patients. This could fail to provide timely healthcare services, which can have serious consequences. Given the critical nature of healthcare IoT applications, rigorous testing at various levels, such as the system level, is essential to ensure their dependability.
@@
Oslo City’s healthcare department [2] collaborates with various healthcare industries to develop IoT applications, aiming to deliver efficient healthcare services to residents. These collaborations typically adhere to DevOps practices, where the Oslo City healthcare team continuously manages, analyzes, and reports problems within the IoT healthcare application under development. One of their activities is to test the REST APIs of such an application at the system level with integrated medical devices and third-party applications. However, manual testing of REST APIs at each evolution stage can be time-consuming and laborious due to the complex and interdependent nature of these APIs, thus necessitating an automated testing solution.
@@
Over time, automated testing of REST APIs has attracted significant interest. Numerous approaches, such as RESTest [3], RESTler [4], Schemathesis [5], RESTCT [6], RestTestGen [7], and EvoMaster [8], have been proposed. The effectiveness of REST API testing approaches has been examined in different empirical studies (e.g., [9, 10, 11]) for various open-source and industrial web applications. Our analysis of the literature revealed two key observations. Firstly, the effectiveness of REST API testing approaches in testing IoT applications’ REST APIs remains unexplored. Secondly, the use of REST API testing approaches in regression testing of evolving REST APIs within the DevOps context is also unexplored. We found only one study by Godefroid et al. [12] that employed RESTler for regression testing of open-source APIs using historical release data. Thus, the evaluation of REST API testing approaches in regression testing of continuously evolving REST APIs of real-world IoT applications remains a research gap. Considering this gap, we assess well-established approaches for testing the REST API of a real-world healthcare IoT application deployed in Oslo City and analyze their practicality in this context.
@@
This paper builds upon our previous work [13] in which we reported an empirical evaluation of RESTest for testing six APIs with 41 endpoints of a healthcare IoT application’s production release. This paper presents a large-scale empirical study that evaluates state-of-the-art and well-established tools—specifically RESTest, EvoMaster, Schemathesis, RESTler, and RestTestGen—for the regression testing of evolving REST APIs associated with a real-world healthcare IoT application, which is progressively developed via the DevOps process. Our experiments encompass 14 different types of releases (daily, weekly, and monthly) and feature all accessible REST APIs of the application under test, i.e., 17 APIs with 120 endpoints. We evaluated REST API testing tools, considering their effectiveness in terms of failures, faults, coverage, regressions, and cost overhead. Results show that EvoMaster generated the highest number of client-side and server-side failures among all the tools. The fault analysis revealed that the tests generated by all tools led to various potential faults, totaling 18 unique faults across 14 releases. The coverage results indicate that RESTest achieved approximately 82% coverage, while EvoMaster, Schemathesis, RESTler, and RestTestGen attained an overall coverage of around 84%. Moreover, analysis of regressions across all 14 releases demonstrated that the tests generated by each tool resulted in the identification of several potential regressions, with EvoMaster at the top. In total, 23 unique regressions were discovered through testing results from all tools across the 14 releases. Through these regressions, we analyzed the stability of APIs during continuous releases, specifically focusing on the state of API implementation (i.e., feature upgrades) and the propagation of failures. The results of this regression analysis offer valuable insights for practitioners, assisting them in identifying potential areas of API vulnerability, determining recurring failures, and choosing a suitable testing tool. Finally, the cost analysis demonstrated that each tool generates a significant number of tests, often leading to numerous duplicate failures, thereby resulting in a cost overhead exceeding 80%.
@@
      Five REST API testing tools, i.e., RESTest, EvoMaster, Schemathesis, RESTler, and RestTestGen configured with different black-box testing techniques, making a total to 10 tools.
@@
      14 rapid releases (including daily, weekly, and monthly) of our industrial healthcare IoT application continuously built in a staging environment following DevOps practices.
@@ 2 Background
We provide an overview of healthcare IoT applications, our industrial context, and the associated challenges.
@@ 2.2 Industrial Context and Challenges
The Directorate of Health in Norway launched the national welfare technology program [14] to offer efficient digital home monitoring and healthcare services. As part of this program, Oslo City’s healthcare department [2] worked with different industry partners to create a cloud-based platform with an integrated network of smart medical devices and healthcare service providers — a platform that resembles the layout of the IoT healthcare applications described in Section 2.1. Oslo City follows the DevOps process to foster these collaborations with industry partners. In this process, Oslo City’s healthcare team continuously manages, analyzes, and reports any issues that arise within the under-development healthcare IoT application. Figure 2 illustrates the commonly followed DevOps process for healthcare IoT applications. In this process, Oslo City initially gets involved during the Plan phase, where they define system requirements and plan various project aspects. The subsequent phases, including Code, Build, Test, Release, and Deploy, involve contributions from industry partners. Upon deploying a specific release, Oslo City’s healthcare team handles the Operate and Monitor phases, in which they analyze the delivered functionalities to ensure dependability and report any issues or necessary enhancements.
@@
Throughout the DevOps process, healthcare IoT applications continuously evolve, adding new services or upgrading existing ones. REST APIs, which make up a significant part of these applications’ backend, also undergo regular evolution. These stages of evolution typically correspond to daily builds for minor changes or bug fixes and weekly builds for major updates or new features. Testing REST APIs at each stage of evolution is labor-intensive and time-consuming. Therefore, automated testing of REST APIs at each evolution stage is essential to ensure the dependability of healthcare IoT applications.
Automated REST API testing of evolving REST APIs presents several challenges. The requirement for frequent testing at each stage of evolution, especially when connected to medical devices and third-party applications, may risk damaging the device or causing service blocking from external applications [1, 15]. Additionally, identifying the root causes of faults or failures due to the dependencies of REST APIs on various third-party applications is a considerable challenge [13]. For testing REST APIs of healthcare IoT applications, generating test data requires specific knowledge of the healthcare domain and an understanding of the application context. Lastly, the complexity and interdependent structure of REST APIs make automated test oracle generation a challenging task. These challenges were primarily addressed in our previous works [15, 16, 17]. In this study, we aim to evaluate REST API testing tools within the context of an evolving healthcare IoT application and to determine the most effective tools for integration into the overall test infrastructure (as indicated in Figure 1).
@@ 3 Related Works
We present related works on testing IoT applications, REST API testing, and empirical evaluations of REST API testing tools.
@@ 3.1 IoT Applications Testing
Numerous studies focus on various aspects of testing IoT applications [18]. These include model-based conformance testing of the IoT [19], user interface-driven acceptance testing of IoT systems [20], testing the heterogeneity of IoT devices in the loop [21], and identifying faults in the integration of devices with IoT applications [22]. Other research areas include combinatorial testing and coverage criteria for IoT systems [23], combinatorial testing for IoT-based smart home applications [24], and simulating the health monitoring activities of healthcare IoT applications [25]. In contrast to the studies mentioned above, our work primarily focuses on the regression testing of evolving REST APIs within a healthcare IoT application. Furthermore, our research involves the empirical evaluation of REST API testing tools, as opposed to proposing a new testing technique. The insights derived from our empirical study could potentially be employed to enhance existing techniques for testing IoT applications across various domains.
@@ 3.2 REST API Testing
Automated testing of REST APIs has gained significant interest, and many approaches have been proposed to address this [26]. These comprise RESTest [3, 27], ARTE [28]—an extension to RESTest, RESTler [4, 29], Schemathesis [5], RESTCT [6], RestTestGen [7, 30], EvoMaster [8, 31], RapiTest [32], QuickREST [33], NLP2REST [34], and ARAT-RL [35]. Some open-source tools like APIFuzzer [36], Tcases [37], and Dredd [38] also support testing of REST APIs. However, as reported in [9], these tools did not outperform research-based tools. In addition, numerous studies target various forms of REST API testing, including model-based testing [39], specification-based testing [40], robustness testing [41], metamorphic testing [42], search-based test case improvement [43], security testing [44], test input validation using deep learning [45], and identifying vulnerabilities in REST APIs [46]. Although these studies propose various approaches for testing REST APIs, our work explicitly evaluates selected tools for testing the REST APIs of a real-world healthcare IoT application within a DevOps environment. In another relevant work, Godefroid et al. [12] proposed an approach for regression testing of REST APIs and evaluated their approach using the RESTler tool on open-source APIs. Compared to this work, our work differs in two aspects. First, our analysis focuses on assessing REST API testing tools in a continuous regression testing environment, whereas their analysis is based on historical release data. Second, we utilized five REST API testing tools along with their supported testing techniques (amounting to a total of 10 tools), in contrast to their work which only considered one tool, specifically RESTler.
@@ 3.3 Empirical Evaluations of REST API Testing Tools
Several studies have conducted empirical research comparing various REST API testing tools. In this regard, Martin et al. [10] compared black-box and white-box testing tools, specifically RESTest and EvoMaster, by using them in four open-source APIs. Corradini et al. [47] also delved into this subject, analyzing four tools: RESTest, RESTler, RestTestGen, and bBOXRT. Their study involved testing the REST APIs of 14 open-source applications. Furthermore, Kim et al. [9] evaluated 10 different tools designed for testing REST APIs, using 20 open-source applications in their study. More recently, Zhang et al. [48] performed an empirical comparison of seven REST API testing tools using 20 different case studies, which included an industrial case study. While the studies outlined above assess various REST API testing tools across different case studies, two key aspects remain unexplored. The first is the effectiveness of these testing tools when applied to REST APIs of large-scale IoT applications. The second is the applicability of these tools for regression testing of evolving REST APIs within a DevOps context. Our work specifically targets these aspects.
In our previous work [13], we evaluated the effectiveness of RESTest in testing six REST APIs of an IoT application in industrial healthcare. In this paper, we extend this scope by evaluating five REST API testing tools for testing continuously evolving REST APIs of a real-world healthcare IoT application in a DevOps context. The specific additions to previous work are outlined below.

### v2

# colorlinks=true, allcolors=black, breaklinks=true

@@ Abstract
Healthcare Internet of Things (IoT) applications often integrate various third-party healthcare applications and medical devices through REST APIs, resulting in complex and interdependent networks of REST APIs. Oslo City’s healthcare department collaborates with various industry partners to develop these applications, enriched with diverse REST APIs that evolve during the DevOps process to accommodate evolving needs such as new features, services, and devices. Oslo City’s primary goal is to utilize automated solutions for continuous testing of REST APIs at each evolution stage to ensure dependability. Although the literature offers various automated REST API testing tools, their effectiveness in regression testing of the evolving REST APIs of healthcare IoT applications within a DevOps context remains undetermined. This paper evaluates state-of-the-art and well-established REST API testing tools—specifically, RESTest, EvoMaster, Schemathesis, RESTler, and RestTestGen—for the regression testing of a real-world healthcare IoT application, considering failures, faults, coverage, regressions, and cost. We conducted experiments using all accessible REST APIs (17 APIs with 120 endpoints), and 14 releases evolved during DevOps. Overall, all tools generated tests leading to several failures, 18 potential faults, up to 84% coverage, and 23 regressions. Over 70% of tests generated by all tools fail to detect failures, resulting in significant overhead.
Keywords Continuous Regression Testing, REST APIs, Web Services, Black-box Testing, Fuzzing, DevOps, CI/CD, Healthcare IoT
@@ 1 Introduction
Healthcare Internet of Things (IoT) applications commonly use a cloud-based architecture as a hub for stakeholders such as healthcare professionals, caretakers, and patients sartaj2023hita . This involves integrating various healthcare applications serving different purposes (such as electronic health records (EHR) for managing patient data) and medical devices (like medicine dispensers) assigned to patients. These integrations into the IoT cloud are facilitated through Application Programming Interfaces (APIs) following the Representational State Transfer (REST) architecture—concisely REST APIs. A malfunction in REST APIs can significantly dismantle the healthcare IoT cloud, potentially making crucial healthcare services unavailable to patients. This could fail to provide timely healthcare services, which can have serious consequences. Given the critical nature of healthcare IoT applications, rigorous testing at various levels, such as the system level, is essential to ensure their dependability.
@@
Oslo City’s healthcare department oslocity collaborates with various healthcare industries to develop IoT applications, aiming to deliver efficient healthcare services to residents. These collaborations typically adhere to DevOps practices, where the Oslo City healthcare team continuously manages, analyzes, and reports problems within the IoT healthcare application under development. One of their activities is to test the REST APIs of such an application at the system level with integrated medical devices and third-party applications. However, manual testing of REST APIs at each evolution stage can be time-consuming and laborious due to the complex and interdependent nature of these APIs, thus necessitating an automated testing solution.
@@
Over time, automated testing of REST APIs has attracted significant interest. Numerous approaches, such as RESTest martin2021restest , RESTler atlidakis2019restler , Schemathesis hatfield2022deriving , RESTCT wu2022combinatorial , RestTestGen corradini2022resttestgen , and EvoMaster arcuri2018evomaster , have been proposed. The effectiveness of REST API testing approaches has been examined in different empirical studies (e.g., kim2022automated ; martin2021black ; arcuri2023building ) for various open-source and industrial web applications. However, REST APIs in IoT-based applications differ significantly due to their hybrid architecture (integrating cloud and edge computing), dependence on heterogeneous devices and third-party services, and frequent vendor-driven API evolution, factors that directly impact their dependability. In addition, healthcare applications, such as the one in our study, enforce stricter integrity, quality, and security requirements due to regulatory compliance, such as the EU’s General Data Protection Regulation (GDPR), requiring more stringent authentication and authorization mechanisms. Our analysis of the literature revealed two key observations. Firstly, the effectiveness of REST API testing approaches in testing IoT applications’ REST APIs remains unexplored. Secondly, the use of REST API testing approaches in the regression testing of evolving REST APIs within the DevOps context is also unexplored. We found only one study by Godefroid et al. godefroid2020differential that employed RESTler for regression testing of open-source APIs using historical release data. Thus, the evaluation of REST API testing approaches in the regression testing of continuously evolving REST APIs of real-world IoT applications remains a research gap. Considering this gap, we assess well-established approaches for testing the REST API of a real-world healthcare IoT application deployed in Oslo City and analyze their practicality in this context.
@@
This paper builds upon our previous work sartaj2023testing in which we reported an empirical evaluation of RESTest for testing six APIs with 41 endpoints of a healthcare IoT application’s production release. This paper presents a large-scale empirical study that evaluates state-of-the-art and well-established tools—specifically RESTest, EvoMaster, Schemathesis, RESTler, and RestTestGen—for the regression testing of evolving REST APIs associated with a real-world healthcare IoT application, which is progressively developed via the DevOps process. Our experiments encompass 14 different types of releases (daily, weekly, and monthly) and feature all accessible REST APIs of the application under test, i.e., 17 APIs with 120 endpoints. We evaluated REST API testing tools, considering their effectiveness in terms of failures, faults, coverage, regressions, and cost overhead. Results show that EvoMaster generated tests led to the highest number of failures among all the tools. The fault analysis revealed that the tests generated by all tools led to various potential faults, totaling 18 unique faults across 14 releases. The coverage results indicate that RESTest achieved approximately 82% coverage, while EvoMaster, Schemathesis, RESTler, and RestTestGen attained an overall coverage of around 84%. Moreover, analysis of regressions across all 14 releases demonstrated that the tests generated by each tool resulted in the identification of several potential regressions, with EvoMaster at the top. In total, 23 unique regressions were discovered through testing results from all tools across the 14 releases. Through these regressions, we analyzed the stability of APIs during continuous releases, specifically focusing on the state of API implementation (i.e., feature upgrades) and the propagation of failures. The results of this regression analysis offer valuable insights for practitioners, assisting them in identifying potential areas of API vulnerability, determining recurring failures, and choosing a suitable testing tool. Finally, the cost analysis demonstrated that each tool generated a significant number of tests, often leading to numerous duplicate failures. For all tools, more than 70% of the generated test cases did not detect any failures, thus incurring considerable overhead.
@@
      Five REST API testing tools, i.e., RESTest, EvoMaster, Schemathesis, RESTler, and RestTestGen configured with different black-box testing techniques, making a total of 10 tools.
@@
      14 rapid releases (including daily, weekly, and monthly) of our industrial healthcare IoT application, continuously built in a staging environment following DevOps practices.
@@ 2 Background
We provide an overview of regression testing, healthcare IoT applications, our industrial context, and the associated challenges.
2.1 Regression Testing
Regressions are unintended defects or behavioral differences between two subsequent versions of the software, typically resulting from various modifications such as bug fixes, the addition of new features, or updates to existing functionality beizer2003software . These regressions often arise from integration issues, especially when modifications in one module impact others due to interdependencies. For instance, a database-related bug fix might inadvertently affect the functionality of the graphical user interface. Similarly, introducing new features or upgrading existing ones may disrupt software behaviors due to potential conflicts with current logic or workflows. Furthermore, regressions may result from the propagation of failures across versions, often triggered by external updates like third-party library or API changes, or the prioritization of critical bugs over lower-priority ones, especially common in rapid release environments. To identify regressions, practitioners use various criteria, including behavioral consistency checks that compare current functionality with expected outcomes from previous versions. Regression testing plays a key role in this regard by systematically running tests on both previous and updated software versions to ensure that recent changes have not adversely impacted existing functionalities yoo2012regression . A straightforward approach to regression testing is the retest-all strategy, which involves rerunning all existing tests on the new version yoo2012regression ; greca2023state . Although simple, this approach can be resource-intensive due to the execution of redundant tests. Therefore, many techniques are also available for selective regression testing, categorized as test case selection, test case minimization, and test case prioritization yoo2012regression . Test case selection techniques focus on choosing a subset of tests that target only the modified software components. Test case minimization techniques aim to remove redundant or obsolete tests from the test suite. Test case prioritization techniques determine the suitable ordering of tests to run with the aim of detecting faults as early as possible. These techniques are particularly useful in Continuous Integration and Continuous Deployment (CI/CD) or DevOps settings, where rapid and frequent releases demand efficient use of testing resources.
@@ 2.3 Industrial Context and Challenges
The Directorate of Health in Norway launched the national welfare technology program wtsoslo to offer efficient digital home monitoring and healthcare services. As part of this program, Oslo City’s healthcare department oslocity worked with different industry partners to create a cloud-based platform with an integrated network of smart medical devices and healthcare service providers—a platform that resembles the layout of the IoT healthcare applications described in Section 2.2. Oslo City follows the DevOps process to foster these collaborations with industry partners. In this process, Oslo City’s healthcare team continuously manages, analyzes, and reports any issues that arise within the under-development healthcare IoT application. Figure 2 illustrates the commonly followed DevOps process for healthcare IoT applications. In this process, Oslo City initially gets involved during the Plan phase, where they define system requirements and plan various project aspects. The subsequent phases, including Code, Build, Test, Release, and Deploy, involve contributions from industry partners. Upon deploying a specific release, Oslo City’s healthcare team handles the Operate and Monitor phases, in which they analyze the delivered functionalities to ensure dependability and report any issues or necessary enhancements.
@@
Throughout the DevOps process, healthcare IoT applications continuously evolve, adding new services or upgrading existing ones. REST APIs, which make up a significant part of these applications’ backend, also undergo regular evolution. These stages of evolution typically correspond to daily builds for minor changes or bug fixes and weekly builds for major updates or new features. Testing REST APIs manually at each stage of evolution is labor-intensive and time-consuming; therefore, automating this testing process is essential to ensure the dependability of healthcare IoT applications. To address this challenge, Oslo City’s primary objective in conducting automated regression testing is to ensure that REST APIs function as intended despite constantly evolving requirements and regular feature upgrades.
Automated REST API testing of evolving REST APIs presents several challenges. The requirement for frequent testing at each stage of evolution, especially when connected to medical devices and third-party applications, may risk damaging the device or causing service blocking from external applications sartaj2023hita ; sartaj2024modelbased . Moreover, identifying the root causes of faults or failures due to the dependencies of REST APIs on various third-party applications is a considerable challenge sartaj2023testing . For testing REST APIs of healthcare IoT applications, generating test data requires specific knowledge of the healthcare domain and an understanding of the application context. Lastly, the complexity and interdependent structure of REST APIs make automated test oracle generation a challenging task. These challenges were primarily addressed in our previous works sartaj2024modelbased ; sartaj2024uncertainty ; sartaj2024medet . In this study, we aim to evaluate REST API testing tools within the context of an evolving healthcare IoT application and to determine the most effective tools for integration into the overall test infrastructure (as indicated in Figure 1).
@@ 3 Related Works
We present related works on testing IoT applications, REST API testing, automated testing in DevOps, and empirical evaluations of REST API testing tools.
@@ 3.1 IoT Applications Testing
Numerous studies focus on various aspects of testing IoT applications dias2018brief . These include model-based conformance testing of the IoT ahmad2016model , user interface-driven acceptance testing of IoT systems leotta2018acceptance , testing the heterogeneity of IoT devices in the loop amalfitano2017towards , and identifying faults in the integration of devices with IoT applications wang2022understanding . Other research areas include combinatorial testing and coverage criteria for IoT systems hu2022ct , combinatorial testing for IoT-based smart home applications garn2022combinatorial , and simulating the health monitoring activities of healthcare IoT applications sotiriadis2014towards . In contrast to the studies mentioned above, our work primarily focuses on the regression testing of evolving REST APIs within a healthcare IoT application. Furthermore, our research involves the empirical evaluation of REST API testing tools, as opposed to proposing a new testing technique. The insights derived from our empirical study could potentially be employed to enhance existing techniques for testing IoT applications across various domains.
@@ 3.2 REST API Testing
Automated testing of REST APIs has gained significant interest, and many approaches have been proposed to address this golmohammadi2022testing . These comprise RESTest martin2021restest ; martin2022online , ARTE alonso2022arte —an extension to—RESTest, RESTler atlidakis2019restler ; godefroid2020intelligent , Schemathesis hatfield2022deriving , RESTCT wu2022combinatorial , RestTestGen corradini2022resttestgen ; corradini2022automated , EvoMaster arcuri2018evomaster ; arcuri2019restful , RapiTest felicio2023rapitest , QuickREST karlsson2020quickrest , NLP2REST kim2023enhancing , ARAT-RL kim2023adaptive , APIRL foley2025apirl , and LlamaRestTest kim2025llamaresttest . Some open-source tools like APIFuzzer apifuzzer , Tcases tcases , and Dredd dredd also support testing of REST APIs. However, as reported in kim2022automated , these tools did not outperform research-based tools. In addition, numerous studies target various forms of REST API testing, including model-based testing liu2022morest , specification-based testing ed2018automatic , robustness testing laranjeiro2021black , metamorphic testing segura2018metamorphic , search-based test case improvement stallenberg2021improving , security testing atlidakis2020checking , test input validation using deep learning mirabella2021deep , and identifying vulnerabilities in REST APIs corradini2023automated . Although these studies propose various approaches for testing REST APIs, our work explicitly evaluates selected tools for testing the REST APIs of a real-world healthcare IoT application within a DevOps environment. In another relevant work, Godefroid et al. godefroid2020differential proposed an approach for regression testing of REST APIs and evaluated their approach using the RESTler tool on open-source APIs. Compared to this work, our work differs in two aspects. First, our analysis focuses on assessing REST API testing tools in a continuous regression testing environment, whereas their analysis is based on historical release data. Second, we utilized five REST API testing tools along with their supported testing techniques (amounting to a total of 10 tools), in contrast to their work, which only considered one tool, specifically RESTler.
3.3 Automated Testing in DevOps
There has been growing interest in developing automated testing approaches for web applications in DevOps across various levels, including unit, integration, and system testing pando2022software ; poth2022integration . Several studies have also focused on specific aspects of testing web applications and services in DevOps, such as security testing rangnau2020continuous , performance-aware regression testing chen2020performance , and reliability testing bertolino2023devopret . However, there is a lack of research on applying automated REST API testing tools for regression testing in DevOps. This gap is particularly significant in the context of IoT-based applications, where REST APIs must operate across hybrid architectures (cloud and edge), accommodate evolving APIs from heterogeneous devices and third-party services, comply with strict data privacy regulations, and maintain consistency across environments such as development, staging, and production. In the context of existing recommended practices for automated testing in DevOps, Abdulla et al. abdulla2023automated provided guidelines for the GitHub development environment, while Patel et al. patel2022state presented general test automation practices within DevOps workflows. However, to the best of our knowledge, we could not find any guidelines specifically addressing the use of REST API testing tools in DevOps for IoT-based healthcare applications, which highlights a unique aspect of our work.
@@ 3.4 Empirical Evaluations of REST API Testing Tools
Several studies have conducted empirical research comparing various REST API testing tools. In this regard, Martin-Lopez et al. martin2021black compared black-box and white-box testing tools, specifically RESTest and EvoMaster, by using them in four open-source APIs. Corradini et al. corradini2021empirical also delved into this subject, analyzing four tools: RESTest, RESTler, RestTestGen, and bBOXRT. Their study involved testing the REST APIs of 14 open-source applications. Furthermore, Kim et al. kim2022automated evaluated 10 different tools designed for testing REST APIs, using 20 open-source applications in their study. More recently, Zhang et al. zhang2023open performed an empirical comparison of seven REST API testing tools using 20 different case studies, which included an industrial case study. While the studies outlined above assess various REST API testing tools across different case studies, two key aspects remain unexplored. The first is the effectiveness of these testing tools when applied to REST APIs of large-scale IoT applications. The second is the applicability of these tools for regression testing of evolving REST APIs within a DevOps context. Our work specifically targets these aspects.
In our previous work sartaj2023testing , we evaluated the effectiveness of RESTest in testing six REST APIs of an IoT application in industrial healthcare. In this paper, we extend this scope by evaluating five REST API testing tools for testing continuously evolving REST APIs of a real-world healthcare IoT application in a DevOps context. The specific additions to previous work are outlined below.